ND-GAr: HPgTPC+ECAL Weekly Meeting

Disappearance Room (WH13SE) and Zoom

Disappearance Room (WH13SE) and Zoom

    • 10:00 AM 10:20 AM
      ND-GAr + WbLS 20m
      Speaker: Michael Wilking (Stony Brook University)

      The ECAL is pressurized, so upstream vs interspersed layers will have similar issues.

      Vivek:  Need a GDML file for the non-uniform ECAL since current geometry in simulation is only .  Also why not put it in SAND?  (Mike notes can't go off.)

      Leo: What is the physics goal?  Mike:  To measure interactions on water-based scintillator if FD is something THEIA-like.  Also thinks that you might need more mass of water.

      Patrick:  T2K FGD isn't in gas volume.  Some concerns about outgassing and materials that won't affect the drift gas.  Alan says that need to pump out volume to vacuum, so water can't leak at 10 bar or vacuum.  So this is tricky.

      Alfons:  Easy way to get a mimimum water target.  But thinks that for a THEIA-like detector might want something that can give more detailed cross section info.  Maybe consider adding a new detector elsewhere, but this limits the off-axis range of PRISM.

      Diego:    Mike is suggested both putting in inside pressure vessel and maybe also another proposal to put inside the TPC volume itself.  Putting it inside the TPC is complicated especially inside the field cage volume.  For putting it inside the ECAL, is 5 cm enough, maybe 10 cm could be reasonable?  More details about proposed thin upstream ECAL needed.  

      Alan: What threshold do you need?  For T2K want to maybe look at protons.  Also a volume between ND-LAr and ND-GAr that could maybe be filled.

      Tom:  If you want good cancellation of systematics, shouldn't it look a bit more like the far detector?  Does this have value without a WbLS far detector?  Maybe to comapre oxygen to argon.

      Diego:  Does a CO2 target help here?  Maybe but you'd need a lot.







    • 10:20 AM 10:40 AM
      Roundtable Updates 20m
      Speakers: Alan Bross (Fermilab), Alfons Weber (JG-Uni Mainz & Fermilab), Diego Gonzalez Diaz, Naseem Khan, Tanaz Mohayai (Fermilab), vivek jain (SUNY Albany)

      GORG: Tanaz: Fixed DAQ issue.  Should be able to take data very soon.

      GOAT: Alan:  Done with IROC tests for now. 

      Light Collection:  Small progress on modelling of optical gain.  Glass GEMs are there and ready to start commissioning at 5 bar next week.

      ECAL: Vivek rediscovered a bug Leo had found in geometry. 

      TOAD:  Electronics plugged in and turned on.  TOAD now taking data with electronics and HV on in air.  So far noise level is well below level expected for signal. OCR is essentially complete and can focus on operating the detector.  Still awaiting magnet ORC.